A rectangular gate 1.2 m wide and 3.5 m long is inclined 35°from a smooth wall. The gate is hinged at point B, and rests against a smooth wall at A which is 2.7 m below the water surface.

a. Compute the force on the gate due to seawater pressure (s.g. = 1.03).

b. Determine the location of the force from the center of gravity of the gate.

c. determine the horizontal force P exerted by the wall at point A.

d. determine the reaction at hinge B.
Answer #1

a w ell 2.7m lo Gate I y=2m 3.5m 350 STDBS 3.5 :.y = 2.00m x=2-87 m csgs • 35 Hinge X = 2.87 m 1 1 F 3.79 3.70 - Вх f 1 c.al© Force P=9 M. . about C-G.(IG) = vertical distance of cp x1-243.53 = 4,2875 m. from free surface :- Yp= ýt IG sinzo А. = 3.7